OpenStudy (anonymous):

What is the value of the discriminant in the equation shown below? x^2+3x-6=0

OpenStudy (anonymous):

Do you have an answer already?

OpenStudy (amistre64):

how is the discriminant defined?

OpenStudy (anonymous):

Given:\[ax^2 + bx + c = 0\]The descriminant is: \[b^2 - 4(a)(c)\] So what do you have for this equation?

OpenStudy (amistre64):

given this explanation of the discriminant; what do you say a,b and c are equal to?

OpenStudy (anonymous):

so it would be 3\[3x^2\]-4(1)(6)

OpenStudy (anonymous):

the descriminant doesn't include the variables.

OpenStudy (anonymous):

i mean no x

OpenStudy (amistre64):

close; you dropped the "-" off the 6 tho

OpenStudy (anonymous):

3^2 - 4(1)(-6) Note that the 6 is negative because it's -6 in the above equation.

OpenStudy (amistre64):

[a] x^2 [+b] x [+c] = 0 [1] x^2 [+3] x [-6] = 0

OpenStudy (anonymous):

so the answer is -15?

OpenStudy (anonymous):

9 - 4(1)(-6) = 9 - (-24) ... \(\ne-15\)

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

oh so 33

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

yay